Sharkey’s Fun Center contains a number of electronic games as well as a miniature golf course and various rides located outside the building. Paul Sharkey, the owner, would like to construct a water slide on one portion of his property. Mr. Sharkey gathered the following information about the slide:

  1. Water slide equipment could be purchased and installed at a cost of $510,000. According to the manufacturer, the slide would be usable for 12 years after which it would have no salvage value.
  2. Mr. Sharkey would use straight-line depreciation on the slide equipment.
  3. To make room for the water slide, several rides would be dismantled and sold. These rides are fully depreciated, but they could be sold for $125,000 to an amusement park in a nearby city.
  4. Mr. Sharkey concluded that about 50,000 more people would use the water slide each year than have been using the rides. The admission price would be $4.90 per person (the same price the Fun Center has been charging for the old rides).
  5. Based on experience at other water slides, Mr. Sharkey estimates that annual incremental operating expenses for the slide would be: salaries, $94,000; insurance, $5,700; utilities, $14,500; and maintenance, $11,300.

Required:

1. Prepare an income statement showing the expected net operating income each year from the water slide.

2-a. Compute the simple rate of return expected from the water slide.

2-b. Based on the above computation, would the water slide be constructed if Mr. Sharkey requires a simple rate of return of at least 12% on all investments?

3-a. Compute the payback period for the water slide.

3-b. If Mr. Sharkey accepts any project with a payback period of five years or less, would the water slide be constructed?
Answer #1
1 Income statement
$ $
Incremental revenue (50000*4.90) 245000
Less:Incremental operating expenses
Salaries 94000
Insurance 5700
Utilities 14500
Maintenance 11300
Depreciation (510000/12) 42500 168000
Net operating income 77000
2-a. Simple rate of return=Net operating income/Initial investment cost=77000/510000=0.15098=15.10%
2-b. Yes.
Since it has a simple rate of return of 15.10% which is more than required return of 12%
3-a. Payback period=Initial investment cost/Annual cash inflow
Annual cash inflow=Net operating income+Depreciation=77000+42500=$ 119500
Payback period=510000/119500=4.27 years
3-b. Yes
Since the payback period of 4.27 years is within the acceptable payback period of 5 years